albinism is an autosomal recessive condition characterized by absence of melanin pigment from the skin, eye and hair. Two carriers of albinism marry and plan to have FOUR children. Assume a 1:1 sex ratio.What is the probability that at least 2 children will be normal?
Answer #1

Gentoype of affected individual = aa

Genotype of normal individual = AA, Aa (carrier)

Aa × Aa

Gametes A a
A AA (normal) Aa (normal, carrier)
a Aa (normal, carrier) aa (affected)

Probability of normal individual = 3/4

Probability of affected individual = 1/4

Probability that atleast 2 children will be normal = 2 children will be normal + 3 children will be normal + 4 children will be normal

Using combinations

4C2 (3/4)^2 (1/4)^2 + 4C3 (3/4)^3 (1/4)^1 + 4C4 (3/4)^4 (1/4)^0

= (4!/2!2!) (3^2)/(4^4) + (4!/3!1!) (3^3)/(4^4) + (4!/4!0!) (3^4)/(4^4)

= (4×3/2)(9)/(256) + (4)(27)/(256) + (1)(81)/(256)

= (63)/256 + (108)/256 + (81)/256

= 252/256

= 0.984

Please rate.